OpenStudy (anonymous):

find a positive number x such that the sum of 16x and 1/x is as small as possible using optimization

OpenStudy (anonymous):

\[f(x)=16x+\frac{1}{x}\]and you want the min right? for \(x\geq 0\)

OpenStudy (anonymous):

then \[f'(x)=16-\frac{1}{x^2}\] find the critical points via \[16-\frac{1}{x^2}=0\] \[\frac{1}{x^2}=16\] \[16x^2=1\] \[x^2=\frac{1}{16}\] \[x=\frac{1}{4}\]

OpenStudy (anonymous):

kind of a long winded way of solving but it is late, you can probably do it shorter. notice i ignored the negative root because you are only concerned with \(x>0\)

OpenStudy (anonymous):

thanks so much